OpenStudy (itrymath):

Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.

OpenStudy (itrymath):

@CamPayne

OpenStudy (itrymath):

@3mar

OpenStudy (itrymath):

it can be a function

OpenStudy (shawn):

All sequences are special type of function.

OpenStudy (itrymath):

but why i dont know

OpenStudy (itrymath):

the numbers dont repeat so this makes it a function

OpenStudy (shawn):

I sequence is basically a function whose domain is the natural number

OpenStudy (itrymath):

so im right?

OpenStudy (shawn):

Well you're right but nor for the reason they you stated. Even if the numbers are repeated, it is still a function.

OpenStudy (itrymath):

WHat how!

OpenStudy (shawn):

consider the sequence, 1,1,1,1,... it is still a function. I.e, f(x) = 1 for all x in the naturals

OpenStudy (itrymath):

then what would make it not a function

OpenStudy (shawn):

ALL sequences are functions. So, there isn't a sequence that is not a function.

OpenStudy (itrymath):

oh okay

OpenStudy (itrymath):

thanks so much m8

OpenStudy (shawn):

no prob

OpenStudy (itrymath):

btw... very well explained !

OpenStudy (mrnood):

A function has only 1 value of 'y' for each value of x in a sequence the 'x' value is the position in the sequence. Ther is only one 'y' value for each position

OpenStudy (mathmate):

By the way, if you're looking for the function of the sequence, it is \(a(i)=\frac{i^2}{2}+\frac{i}{2}+5\) for i\(\ge\)0 which is, as discussed above, a function.

OpenStudy (itrymath):

wooooow @mathmate dont pull some complicate stuff on me (rofl)

OpenStudy (mathmate):

@ItryMath Actually, I forgot to add: "all polynomials are functions". That can be useful for certain questions.

OpenStudy (itrymath):

thanks @mathmate

OpenStudy (mathmate):

You're welcome! :)